# Controllable Patching for Compute-Adaptive Surrogate Modeling of Partial Differential Equations

### v1

@@ Abstract
Patch-based transformer surrogates have become increasingly effective for modeling spatiotemporal dynamics, but the fixed patch size is a major limitation for budget-conscience deployment in production. We introduce two lightweight, architecture-agnostic modules—the Convolutional Kernel Modulator (CKM) and Convolutional Stride Modulator (CSM)—that enable dynamic patch size control at inference in patch-based models, without retraining or accuracy loss. Combined with a cyclic patch-size rollout, our method mitigates patch artifacts and improves long-term stability for video-like prediction tasks. Applied to a range of challenging 2D and 3D PDE benchmarks, our approach improves rollout fidelity and runtime efficiency. To our knowledge, this is the first framework to enable inference-time patch-size tunability in patch-based PDE surrogates. Its plug-and-play design makes it broadly applicable across architectures—establishing a general foundation for compute-adaptive modeling in PDE surrogate tasks  [^1] Full codebase will be made publicly available upon publication in this repository..
@@ 1 Introduction
Numerical methods have long been the gold standard for simulating spatio-temporal dynamics in systems governed by partial differential equations (PDEs). Numerical solvers offer convergence guarantees, exploit problem sparsity, and provide fine-grained control of the accuracy-compute trade-off through configurable resolution [1, 2]. This last aspect is of vital importance as it is well-known that certain aspects of physics can only be resolved with sufficient resolution so that the choice of resolution is driven by both physical necessity and the availability of compute[3, 4].
Figure 1: Overview diagram of CKM and CSM approaches for adaptive patching in ViT-based PDE surrogate models. Here, $k$ and $s$ denote the convolutional kernel size and stride, respectively, which control the downsampling rate. Additionally, this figure highlights our rollout strategy at inference, where the dynamic kernel or stride alternates cyclically between 4, 8, and 16 at each timestep.
In recent years, deep learning and particularly transformers have emerged as a powerful framework for training surrogate models [5, 6, 7, 8]. Often derived from computer vision methods such as the Vision Transformer (ViT [9]), these models segment discretized fields into distinct, non-overlapping patches, converting them into tokens for the transformer. This results in a downsampling of the input, an essential step as self-attention cost scales quadratically with token count. While initially expensive to train, this cost can be amortized by repeated usage in problems such as forecasting [10], PDE-constrained optimization [11], and parameter inference [12, 13].
However, the inflexible computational cost of the ViT-based framework can become an undue burden as models are deployed in production. At a fixed input resolution, a ViT-based model incurs a fixed cost, regardless of problem complexity or the budget of the user. This problem is not unique to physical simulation - notably, NLP researchers have explored variable-cost solutions, such as switching between models of different expense [14], adjusting hidden dimension size [15], or other test-time scaling methods [16].
While the problem is not unique to physical settings, the physics domain does offer more natural paths toward adaptive compute. It is well understood in physical modeling that there are often hard resolution thresholds required for resolving fine-scale features like shocks or wave-fronts [17]. In many problems, these small scales can influence larger scales overtime and thus resolving them can lead to significant accuracy gains. For ViT-style models, the internal resolution is entirely determined by the patch size used for tokenization. This can influence both model accuracy and computational cost [9, 18]. Smaller patches act to reduce the level of downsampling, allowing the transformer to resolve finer details with greater accuracy (Fig. 3) but at significant cost. Thus, we seek a compute-adaptive approach for balancing the accuracy and budget demands of users by controlling the model’s internal patch resolution.
To address this issue, this paper introduces two different strategies for adjusting downsampling rates in patch-based tokenization blocks enabling controllable resolution transformers and giving users more precise control over their desired location on compute-accuracy frontier. Convolutional Kernel Modulation (CKM) utilizes the patch-resizing of FlexiViT [19] in both the encoder and decoder while Convolutional Stride Modulation (CSM) utilizes a static convolutional kernel but adjusts the compression rate directly by modulating the stride. These two strategies are visualized in Fig. 1.
Crucially, this flexibility opens up new modeling capabilities that are simply unavailable to static-patch architectures. For example, we show that during autoregressive rollout — where errors can accumulate over time — alternating patch sizes across timesteps mitigates patch-induced artifacts such as harmonic spectral spikes (see Fig. 2). This artifact reduction highlights the emergent benefits of enabling patch/stride diversity at inference, leading to broader implications for stability and robustness in long-horizon surrogate modeling.
Figure 2: Left: Residual power spectrum of the density field at the 20th rollout step for CKM, CSM, and fixed patch size 16 (p16) on the turbulent_radiative_layer_2D dataset [20]. Harmonic spikes indicate patch-induced artifacts, significantly reduced in CKM and CSM. Right: Prediction at rollout step 20, where CKM and CSM mitigate visible patching effects in fixed patch 16 (P16) model.
@@
We propose a general framework for enabling compute-adaptive inference in transformer-based PDE surrogates. Our contributions span model design, architectural generality, and empirical insights:
@@
    Controllable Tokenization: We introduce Convolutional Kernel Modulation (CKM) and Convolutional Stride Modulation (CSM) as two approaches for controllable tokenization. Used with randomized compression rates during training, they allow test-time compute-adaptive accuracy.
@@
    Architecture-agnostic deployment: These flexible patching and striding strategies operate at the tokenization level of the encoder and decoder, and are compatible with a broad class of transformer architectures. We show this generality using two different types of standard base transformer processors – vanilla ViT [9] with full attention and axial ViT [21].
@@
    Competitive to fixed patch size baselines: Across a diverse suite of 2D and 3D PDE datasets from the Well collection [20], CKM and CSM consistently match or outperform fixed-patch size surrogates at multiple token counts—despite being trained once. This holds for both next-step prediction and long autoregressive rollouts. Our models are also shown to be competitive with a range of state-of-the-art non-patch-based baselines, underscoring their broader importance.
@@
    Patch Artifact Reduction: Leveraging the CKM and CSM models mentioned above, we introduce a novel rollout strategy that mitigates patch-induced artifacts (spikes at harmonics of the inverse patch size, see Fig. 2). This significantly enhances the quality of long-term predictions, especially in capturing finer-scale physical features. This artifact reduction emerges naturally from the ability to modulate tokenization over time—demonstrating an additional benefit of flexible inference.
2 Background and related work
Notations. We focus on autoregressive surrogate models for time-evolving PDE systems, where the goal is to predict the next state of a discretized physical field from a short temporal context. This setup mirrors video prediction on a uniform grid. We refer to our proposed CKM/CSM models as compute-elastic, meaning they support tunable patch or stride resolution at inference time. Unless otherwise stated, scalability refers to this token-level flexibility, not mesh-level resolution changes.
@@ 2 Background and related work
Architecture-agnostic tokenization for PDE surrogates. CKM and CSM are architecture-agnostic tokenization strategies applied at the encoder and decoder stages. They control the patch/stride configuration for spatial downsampling while remaining independent of the transformer architecture. We demonstrate compatibility with standard vanilla ViTs (full attention) [9] and axial ViTs [22]. This modularity makes CKM and CSM flexible plug-ins for a wide range of transformer-based PDE surrogates, enabling adaptive deployment without redesign. To our knowledge, this is the first work to propose such flexible, architecture-agnostic tokenization for PDE surrogates.
@@
Surrogate modeling of PDEs with deep learning. PINNs [23, 24, 25, 26, 27, 28, 29, 30, 31] embed physics in loss functions, while CNN- and GNN-based surrogates operate on structured or unstructured meshes, supporting spatiotemporal tasks such as climate and fluid modeling [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Neural operators [47, 48, 49, 50, 51, 52, 53, 5, 54, 55, 56, 57, 58] aim for resolution-invariant learning, with FNOs [59] using Fourier transforms for global context.
@@
ViTs and compute adaptive models for PDE surrogate modeling. Initially developed for image recognition tasks, ViTs [9] have seen increasing adoption in the PDE surrogate modeling literature [60, 6, 61]. Despite their strengths, traditional ViTs rely on fixed patch sizes, which strongly limits their flexibility and adaptability when modeling PDE–based systems as discussed in the introduction. Adaptive patching enables dynamic adjustment of patch sizes based on task complexity, as seen in FlexiViT [19] for vision tasks, which adapts ViTs to varying patch sizes and compute budgets. Although explored in computer vision, adaptive patching for PDE surrogate modeling has, to our knowledge, only been explored in [62], the difference with our work being that their patchification is data-driven, while ours can be manually controlled based on compute requirements.
3 Flexible encoder & decoder frameworks
@@
Overview. CKM and CSM operate as encoder-decoder modules within a ViT based architecture for autoregressive PDE prediction. The transformer core applies a sequence of temporal attention, spatial attention, and MLP blocks (App. B.2). Crucially, our framework is architecture-agnostic—we test it with axial attention and vanilla ViTs (and also a recent model– Continuous ViT [63] in App. D).
@@
Utility of compute-elastic models. As illustrated in Fig.3, increasing the number of tokens—equivalent to using smaller patch sizes in models with fixed patches—enhances the accuracy of autoregressive prediction tasks, but at a higher computational cost. The cost for models trained with fixed patch sizes cannot be controlled at inference time. This limits the utility of these models in budget-aware production settings and motivate the exploration of compute-elastic models for PDE modeling with ViT-inspired architectures.
@@
Research in the broader machine learning community has extensively demonstrated the value of scale [64, 65]. As transformer-based surrogate models similarly scale in parameter count, ensuring inference-time scalability is crucial, particularly to ensure these models are able to be used in compute-constrained environments. To avoid any confusion, by inference scalability, we mean patching/striding scalability or tunability, and not scalability in mesh resolution. Training and maintaining multiple fixed-patch models is a burden especially as very large models, such as large pretrained or foundation models are being built [7, 8, 66]. In such a scenario, if our compute-elastic strategies can enable training of a single model without retraining for multiple patch resolutions without any loss in accuracy, these strategies can form the building blocks for larger scale foundation models for PDEs.
@@
Given an input image with height $H$ and width $W$, ViTs will divide the image into non-overlapping patches of size $k\times k$, producing $N=\lfloor H/k\rfloor\cdot\lfloor W/k\rfloor$ tokens. For global attention, transformer cost scales as $\mathcal{O}(N^{2})$, and for axial attention, as $\mathcal{O}(N\sqrt{N})$ [22]. In convolutional tokenization, patches are extracted via kernels of size $k$ and stride $s$, resulting in a token count
$N\coloneqq N_{h}\cdot N_{w},\quad\text{for}\quad N_{h}\coloneqq\left\lfloor\frac{H-k}{s}\right\rfloor+1,\quad N_{w}\coloneqq\left\lfloor\frac{W-k}{s}\right\rfloor+1,\quad$ (1)
While most ViT based architectures use $k=s$, our CKM and CSM modules decouple these parameters to flexibly control token count at inference.
@@
CKM introduces dynamic patch size selection within convolutional encoder and decoder blocks, enabling a model to tokenize input fields using patch sizes $k$in {4, 8, 16} at each forward pass. These powers-of-two sizes reflect standard discretizations in PDE data, and contrast with most ViT-based surrogates that fix patch size at 16 throughout training and inference [7, 66].
Consider input $x\in\mathcal{X}^{T}$ for some input space $\mathcal{X}^{T}\subseteq\mathbb{R}^{M\times H\times W\times T\times C}$, with $M,T$ and $C$ being the batch size, time context and input channels respectively. CKM resizes a base convolutional kernel $w^{\text{base}}\in\mathbb{R}^{k_{\text{base}}\times k_{\text{base}}\times C\times C^{\prime}}$ using PI–resize [19] (App. A):
@@
$w=B^{\mathsf{T}\dagger}w^{\text{base}},\quad x_{\text{enc}}=\operatorname{Conv}(x,w,\texttt{stride}=k),$ (2)
where $B\in\mathbb{R}^{k\times k_{\text{base}}}$ is a bicubic interpolation matrix and $B^{\mathsf{T}\dagger}$ its pseudoinverse. Tokens are processed by the transformer processor (this transformer can have any architecture), producing latents $x_{\text{lat}}\in\mathcal{X}_{\text{lat}}$ for some latent space $\mathcal{X}_{\text{lat}}$. The decoder reuses the same $k$ and $w$ to reconstruct:
$\hat{x}=\operatorname{ConvTranspose}(x_{\text{lat}},w,\texttt{stride}=k),$ (3)
yielding a next-step prediction $\hat{x}\in\mathcal{X}^{1}\subseteq\mathbb{R}^{B\times H\times W\times 1\times C}$. Pseudocode is in App. A.5.
Comparison with FlexiViT. While FlexiViT introduced flexible patching in image classification via a single-stage encoder [19], it lacked a decoder and was not designed for spatiotemporal forecasting. CKM generalizes this idea, enabling PI–resize modulation across any convolutional stage—encoder or decoder, single- or multi-stage—as long as convolutional weights are available. FlexiViT is thus a special case within the broader CKM framework.
Modular Applicability. In our vanilla ViT and Axial ViT models, we adopt two-stage convolutional encoders and decoders following the hMLP architecture [67], which has recently been adopted for large-scale PDE surrogates [66, 7], and apply CKM independently at each stage.
Convolutional Stride Modulator (CSM).
CSM enables flexible downsampling by sampling the stride $s\in\{4,8,16\}$ at each forward pass. Given input $x\in\mathcal{X}^{T}$, we pad the input with learned tokens (informed by boundary conditions) to avoid edge artifacts. Using a fixed kernel $k^{\text{base}}$, the encoder produces:
$x_{\text{enc}}=\operatorname{Conv}(x,k^{\text{base}},\texttt{stride}=s),$ (4)
which is passed through a transformer to obtain $x_{\text{lat}}$. The decoder applies a transposed convolution:
$\hat{x}=\operatorname{ConvTranspose}(x_{\text{lat}},k^{\text{base}},\texttt{stride}=s),$ (5)
to generate the autoregressive prediction $\hat{x}\in\mathcal{X}^{1}$. Like CKM, CSM is compatible with both single-stage and multi-stage patching schemes. Together, CKM and CSM offer general strategies for flexifying convolutional tokenization in ViTs. Pseudocode is in App. A.6.
Novel rollout strategy.
In traditional autoregressive forecasting, patch size is fixed across time. Our compute-elastic modules unlock a novel strategy: alternating patching during rollout. At each step, the model cycles through $k\in\{4,8,16\}$ for CKM (or $s\in\{4,8,16\}$ for CSM). This test-time scheduling is entirely decoupled from training and enables new temporal patterns in tokenization.

### v2

# Overtone: Cyclic Patch Modulation for Clean, Efficient, and Flexible Physics Emulators

@@ Abstract
Transformer-based PDE surrogates achieve remarkable performance but face two key challenges: fixed patch sizes cause systematic error accumulation at harmonic frequencies, and computational costs remain inflexible regardless of problem complexity or available resources. We introduce Overtone, a unified solution through dynamic patch size control at inference. Overtone’s key insight is that cyclically modulating patch sizes during autoregressive rollouts distributes errors across the frequency spectrum, mitigating the systematic harmonic artifact accumulation that plague fixed-patch models. We implement this through two architecture-agnostic modules—CSM (using dynamic stride modulation) and CKM (using dynamic kernel resizing)—that together provide both harmonic mitigation and compute-adaptive deployment. This flexible tokenization lets users trade accuracy for speed dynamically based on computational constraints, and the cyclic rollout strategy yields up to 40% lower long rollout error in variance-normalised RMSE (VRMSE) compared to conventional, static-patch surrogates. Across challenging 2D and 3D PDE benchmarks, one Overtone model matches or exceeds fixed-patch baselines across inference compute budgets, when trained under a fixed total training budget setting. Code is made available at: https://github.com/payelmuk150/patch-modulator.
@@ 1 Introduction
Numerical solvers have long been the gold standard for simulating spatio-temporal dynamics in systems governed by partial differential equations (PDEs). Numerical solvers offer convergence guarantees, exploit problem sparsity, and provide fine-grained control of the accuracy-compute trade-off through configurable resolution (Morton and Mayers, 2005; Malalasekera, 2007), driven by both scientific necessity and computational availability (Berger and Colella, 1989; Wedi, 2014).
Deep learning has established itself as a powerful framework for training surrogate models of physical simulations (Cao, 2021; Li et al., 2023; McCabe et al., 2023; Herde et al., 2024). Though expensive to train, relative cheap inference costs result in overall compute savings, which has led deep learning surrogates to become popular in forecasting (Bi et al., 2023), PDE-constrained optimization (Li et al., 2022), and parameter inference (Cranmer et al., 2020; Lemos et al., 2023). Recent approaches are often derived from computer vision methods such as Vision Transformers (ViT (Dosovitskiy et al., 2020)), which segment discretized fields into patches and then tokens. Non-overlapping patches are used as a means to reduce token count and the associated quadratic attention cost.
However, these surrogates face two major limitations. First, we discover that in autoregressive rollouts, fixed patch sizes cause systematic error accumulation at harmonic frequencies through temporal coherence. While patch artifacts themselves are well-known, we show that repeatedly hitting the same harmonics at every timestep creates a distinct accumulation pattern. When patches of size $k$ are used consistently during autoregressive rollouts, the artificial boundaries inject errors at wavenumbers $n/k$ where $n$ is an integer. These errors constructively interfere over time, creating pronounced spectral spikes (Figure 2) and visible grid-like artifacts. Second, fixed-patch models have inflexible computational cost. In physical modeling, different applications have hard resolution thresholds for resolving features like shocks or wave-fronts (Boyd, 2001), yet smaller patches that improve accuracy (Dosovitskiy et al., 2020; Wang et al., 2025a) cannot be selected post-training.
Figure 1: Illustration of Overtone’s CSM (stride modulation) and CKM (kernel modulation). Cyclic modulation distributes errors across frequencies, preventing error accumulation at harmonics and enabling accuracy-compute trade-offs at inference time.
We address both challenges through dynamic patch size control during rollout—so far unexplored in autoregressive modeling. Cyclically alternating patch sizes between $k_{1}$, $k_{2}$, and $k_{3}$ at consecutive timesteps can mitigate artifacts from accumulating coherently at single harmonics, instead distributing them across the frequency spectrum. This same capability enables compute-adaptive deployment without retraining, allowing users to balance accuracy and computational demands based on available resources.
Overtone implements this through two modules: Convolutional Stride Modulation (CSM) maintains a static kernel but dynamically modulates stride, while Convolutional Kernel Modulation (CKM) uses bicubic interpolation to resize a single kernel (as in (Xu et al., 2014; Beyer et al., 2023)) for both encoder and decoder, all while cyclically modulating these during rollout (Figure 1). We empirically demonstrate that this cyclic modulation fundamentally alters artifact accumulation in autoregressive rollouts, also with motivation of the technique in spectral analysis (Figure 2), to show how temporal coherence in tokenization has been an unaddressed source of error in spatiotemporal models.
Figure 2: Left: Averaged 1D residual power spectrum at rollout step 20 revealing harmonic error accumulation in fixed-patch models. The fixed patch 16 model shows pronounced spikes at harmonic frequencies $k/16$, while CSM and CKM with cyclic modulation distribute spectral errors across the frequency range, supporting our theoretical motivation (Section 2). Note that this is residual spectrum wrt the ground truth, which is trivially zero for the true simulation, and is therefore not shown. Right: Spatial manifestation of harmonic artifacts at rollout step 20 on Turbulent Radiative Layer 2D. Cyclic patch modulation eliminates the grid-like distortions in the fixed model, showing the practical impact of Overtone’s rollout strategy.
@@
We present Overtone, a framework that identifies and addresses a fundamental issue in patch-based surrogates while providing practical compute flexibility:
@@
    Harmonic Artifact Diagnosis: We identify that fixed-patch tokenization causes systematic spectral error accumulation at harmonic frequencies $k/p$, manifesting as spectral spikes and spatial artifacts that degrade long-term predictions in autoregressive surrogates.
@@
    Cyclic Rollout Strategy: We introduce cyclic stride/patch modulation during inference for autoregressive rollouts. While adaptive patching exists in vision, we are not aware of prior work that cyclically modulates patches during autoregressive rollouts in computer vision, video modeling, or PDE applications. The rollout cycles tokenization scale (e.g., 4 $\rightarrow$ 8 $\rightarrow$ 16, repeated), which (i) prevents errors frequency space errors from repeatedly reinforcing at a single patch scale—thereby reducing patch-lattice harmonic artifacts—and (ii) periodically leverages finer tokenization (4 and 8) for higher-fidelity steps while retaining the efficiency of coarser patches (16). Together, these effects reduce VRMSE by up to 40% relative to fixed-patch baselines (patch size 16), without retraining.
@@
    Architecture Agnostic Controllable Tokenization Modules: We develop CSM (using dynamic stride adjustment) and CKM (using kernel interpolation (Xu et al., 2014; Beyer et al., 2023)) as architecture-agnostic methods for controllable tokenization. These modules operate at the tokenization level and are compatible with backbones like vanilla ViT Dosovitskiy et al. (2020), axial ViT Ho et al. (2020), CViT Wang et al. (2025b), making them flexible plug-ins for PDE surrogates.
@@
    Compute-Adaptive Deployment: Beyond strong accuracy gains, Overtone provides practical trade-offs between speed and precision at inference. A single model can adjust patch sizes based on available resources, matching or exceeding multiple fixed-patch baselines across all compute budgets on challenging 2D/3D PDE benchmarks from the Well (Ohana et al., 2024). Overtone is also competitive with a range of SOTA non-patch baselines, underscoring its broader significance.
We focus on autoregressive surrogate models for time-evolving PDE systems, where the goal is to predict the next state of a discretized physical field from a short temporal context.
1.1 Background and related work
Architecture-agnostic tokenization for PDE surrogates.
@@ 1.1 Background and related work
CSM and CKM are architecture-agnostic tokenization strategies applied at the encoder and decoder stages. They control the patch/stride configuration for spatial downsampling while remaining independent of the transformer architecture. We show compatibility with standard vanilla ViTs (full attention) (Dosovitskiy et al., 2020) and axial ViTs (Ho et al., 2019). This modularity makes CSM and CKM flexible plug-ins for diverse transformer-based PDE surrogates, allowing adaptive deployment without redesign.
Surrogate modeling of PDEs with deep learning.
@@
PINNs (Raissi et al., 2019; Cai et al., 2021; Han et al., 2018; Hao et al., 2023a; Karniadakis et al., 2021; Lu et al., 2021; Pang et al., 2019; Sun et al., 2020; Zhu et al., 2023) embed physics in loss functions, while CNN- and GNN-based surrogates operate on structured or unstructured meshes, supporting spatiotemporal tasks such as climate and fluid modeling (Brandstetter et al., 2022c; Gupta and Brandstetter, 2022; Janny et al., 2023; Li and Farimani, 2022; Lötzsch et al., 2022; Pfaff et al., 2021; Prantl et al., 2022; Sanchez-Gonzalez et al., 2020; Stachenfeld et al., 2022; Thuerey et al., 2020; Ummenhofer et al., 2020; Wang et al., 2020; Lam et al., 2022; Nguyen et al., 2023; Pathak et al., 2022). Neural operators (Li et al., 2020b; c; Lu et al., 2019; 2022; Ovadia et al., 2023; Brandstetter et al., 2022a; b; Cao, 2021; Gupta et al., 2021; Hao et al., 2023b; Jin et al., 2022; Kissas et al., 2022; Kovachki et al., 2021) aim for resolution-invariant learning, with FNOs (Li et al., 2020a) using Fourier transforms for global context.
ViTs and compute adaptive models for PDE surrogate modeling.
@@
Initially developed for image recognition tasks, ViTs (Dosovitskiy et al., 2020) have seen increasing adoption in the PDE surrogate modeling literature (Cao, 2021; Li et al., 2023; Liu et al., 2023). Despite their strengths, traditional ViTs rely on fixed patch sizes, which strongly limits their flexibility and adaptability when modeling PDE-based systems as discussed in the introduction. Examples of kernel interpolation for adaptive sizing include SiCNN (Xu et al., 2014), which applied bicubic interpolation to CNN kernels and trained at multiple scales, and FlexiViT (Beyer et al., 2023), which applies bicubic interpolation to ViT kernels for image classification. Although explored in computer vision, adaptive patching for PDE surrogate modeling has only been explored in (Zhang et al., 2024a), the difference with our work being that their patchification is data-driven, while ours can be manually controlled based on compute requirements.
2 Methods
Overview.
@@ 2 Methods
CSM and CKM operate as encoder-decoder modules within a ViT based architecture for autoregressive PDE prediction. The transformer core applies a sequence of temporal attention, spatial attention, and MLP blocks (Section C.2). Crucially, Overtone’s framework is architecture-agnostic—we test it with axial attention and vanilla ViTs (and also a recent model—Continuous ViT (Wang et al., 2025b) in Appendix E).
Utility of flexible models.
@@
As illustrated in Figure 3, increasing the number of tokens—equivalent to using smaller patch sizes in models with fixed patches—enhances the accuracy of autoregressive prediction tasks, but at a higher computational cost. The cost for models trained with fixed patch sizes cannot be controlled at inference time. This limits the utility of these models in budget-aware production settings and motivate the exploration of flexible models for PDE modeling with ViT-inspired architectures.
@@
Research in the broader machine learning community has extensively demonstrated the value of scale (Zhai et al., 2022; Hoffmann et al., 2022). As transformer-based surrogate models similarly scale in parameter count, ensuring inference-time scalability is crucial, particularly to ensure these models are able to be used in compute-constrained environments. To avoid any confusion, by inference scalability, we mean patching/striding scalability or tunability, and not scalability in mesh resolution. Training and maintaining multiple fixed-patch models is a burden especially as large pretrained or foundation models are being built (McCabe et al., 2023; Herde et al., 2024; Morel et al., 2025). In such a scenario, if Overtone’s flexible strategies can enable training of a single model without retraining for multiple patch resolutions without any loss in accuracy, these strategies can form the building blocks for larger scale foundation models for PDEs.
@@
Given an input image with height $H$ and width $W$, ViTs will divide the image into non-overlapping patches of size $k\times k$, producing $N=\lfloor H/k\rfloor\cdot\lfloor W/k\rfloor$ tokens. For global attention, transformer cost scales as $\mathcal{O}(N^{2})$, and for axial attention, as $\mathcal{O}(N\sqrt{N})$ (Ho et al., 2019). In convolutional tokenization, patches are extracted via kernels of size $k$ and stride $s$, resulting in token count $N=N_{h}\cdot N_{w}$ where $N_{h}=\lfloor(H-k)/s\rfloor+1$ and $N_{w}=\lfloor(W-k)/s\rfloor+1$. While most ViT based architectures use $k=s$, our CSM and CKM modules decouple these parameters to flexibly control token count at inference.
Convolutional Stride Modulator (CSM).
CSM provides flexible downsampling by modulating the stride $s\in\{4,8,16\}$ at each forward pass. Consider input $x\in\mathcal{X}^{T}$ for some input space $\mathcal{X}^{T}\subseteq\mathbb{R}^{M\times H\times W\times T\times C}$, with $M,T$ and $C$ being the batch size, time context and input channels respectively. We pad the input with learned tokens (informed by boundary conditions) to avoid edge artifacts. During training, the model is trained with random stride sizes sampled from 4, 8, 16 during the forward pass (Section K.5). At inference, starting from input context $\hat{x}^{0}=(x^{1},...,x^{T})$, CSM uses a fixed kernel $w^{\text{base}}$ and computes:
$x_{\text{enc}}^{i}=\mathop{\mathrm{Conv}}_{\text{stride }s_{i}}(\hat{x}^{i},w^{\text{base}}),\quad x_{\text{lat}}^{i}=\mathop{\mathrm{Processor}}(x_{\text{enc}}^{i}),\quad\hat{x}^{i+1}_{T}={\mathop{\mathrm{Conv}}_{\text{stride }s_{i}}}^{\mathsf{T}}(x_{\text{lat}}^{i},w^{\text{base}}),$ $\text{where we cycle }s_{i}=(4,8,16)_{(i\bmod 3)+1}\text{ and slide }\hat{x}^{i+1}_{1:T-1}=\hat{x}^{i}_{2:T}.$
Here, $\mathop{\mathrm{Processor}}$ denotes a transformer-based processor that operates on the encoded tokens, ${\mathop{\mathrm{Conv}}}^{\mathsf{T}}$ indicates a transposed convolution, and the output $\hat{x}^{i+1}_{T}\in\mathcal{X}^{1}\subseteq\mathbb{R}^{B\times H\times W\times 1\times C}$ is the next-step prediction. CSM is compatible with both single-stage and multi-stage encoder-decoder modules. Implementation details and code are provided in the supplementary materials.
@@
CKM introduces dynamic patch size selection within convolutional encoder and decoder blocks, enabling models to select a patch size $k\in\{4,8,16\}$ for each forward pass—powers-of-two that reflect standard discretizations in PDE data. Unlike most ViT-based surrogates that fix patch size at 16 (McCabe et al., 2023; Morel et al., 2025), CKM provides flexible tokenization essential for harmonic error mitigation (analyzed in Section 2). This is achieved using kernel interpolation (Xu et al., 2014; Beyer et al., 2023).
At training time, the model sees a random kernel size sampled from 4, 8, 16 (Section K.6). At inference, starting from $\hat{x}^{0}=(x^{1},...,x^{T})$, CKM resizes a base convolutional kernel $w^{\text{base}}\in\mathbb{R}^{k_{\text{base}}\times k_{\text{base}}\times C\times C^{\prime}}$ using PI-resize (detailed in Appendix K) and computes:
@@
$x_{\text{enc}}^{i}=\mathop{\mathrm{Conv}}_{\text{stride }k_{i}}(\hat{x}^{i},B_{k_{i}}^{\mathsf{T}\dagger}w^{\text{base}}),\quad x_{\text{lat}}^{i}=\mathop{\mathrm{Processor}}(x_{\text{enc}}^{i}),\quad\hat{x}^{i+1}_{T}={\mathop{\mathrm{Conv}}_{\text{stride }k_{i}}}^{\mathsf{T}}(x_{\text{lat}}^{i},B_{k_{i}}^{\mathsf{T}\dagger}w^{\text{base}}),$ $\text{where we cycle }k_{i}=(4,8,16)_{(i\bmod 3)+1}\text{ and slide }\hat{x}^{i+1}_{1:T-1}=\hat{x}^{i}_{2:T}.$
Here, $B_{k_{i}}\in\mathbb{R}^{k_{i}\times k_{\text{base}}}$ is a bicubic interpolation matrix and $B_{k_{i}}^{\mathsf{T}\dagger}$ its pseudoinverse transpose.
Examples of kernel interpolation for adaptive sizing include SiCNN (Xu et al., 2014), which applied bicubic interpolation to CNN kernels and trained at multiple scales, and FlexiViT (Beyer et al., 2023), which applies bicubic interpolation to ViT kernels for image classification. Using kernel interpolation, we introduce cyclic modulation during rollouts—transforming it from a compute-flexibility tool into a method for mitigating harmonic error accumulation in long-horizon predictions. In our vanilla ViT and Axial ViT models, we adopt two-stage convolutional encoders and decoders following the hMLP architecture (Touvron et al., 2022), which has recently been adopted for large-scale PDE surrogates (Morel et al., 2025; McCabe et al., 2023), and apply CKM independently at each stage.
Together, the Overtone framework offers general strategies for flexifying convolutional tokenization in ViT-based autoregressive schemes, as well as mitigating accumulating patch artifacts.
Cyclic rollout strategy.
In standard autoregressive forecasting, patch sizes remain fixed across the rollout. By contrast, Overtone’s flexible modules enable patch sizes or strides to alternate across timesteps, introducing new temporal patterns in tokenization. This test-time capability, decoupled from training, leads to a striking empirical effect: alternating rollouts consistently suppress patch artifacts—periodic errors that appear as harmonics in spectral residuals (Figure 2)